### v3

# Conformal Killing gravity in static spherically-symmetric spacetimes

Abstract
We identify an anisotropic divergence-free conformal Killing tensor $K_{jl}$ for static spherically symmetric spacetimes, and write the conformal Killing gravity equations as Einstein equations augmented by this tensor. The field equations are of second order: this fact allows for analytic solutions and considerably simplifies the derivation of results of previous studies based on the original Harada equations. In particular, we prove the equivalence of the known third order field equations, with the second order ones obtained by us in the conformal Killing parametrization.
The structure of the Ricci tensor and of the conformal Killing tensor are compatible with both anisotropic fluid sources and (non)-linear electrodynamics. We reobtain covariantly and in simple steps the general static spherical solutions for vacuum and linear electrodynamics. Moreover we recover the purely magnetic Lagrangian functions that induce metrics of interest for black holes.
I Conformal Killing gravity
Recently J. Harada [1, 2] introduced a new theory of gravity, with field equations
$H_{jkl}=$ $8\pi\,T_{jkl}$ (1)
$H_{jkl}=$ $\nabla_{j}R_{kl}+\nabla_{k}R_{lj}+\nabla_{l}R_{jk}$
$-\tfrac{1}{3}(g_{kl}\nabla_{j}R+g_{lj}\nabla_{k}R+g_{jk}\nabla_{l}R)$
$T_{jkl}=$ $\nabla_{j}T_{kl}+\nabla_{k}T_{lj}+\nabla_{l}T_{jk}$
$-\tfrac{1}{6}(g_{kl}\nabla_{j}T+g_{lj}\nabla_{k}T+g_{jk}\nabla_{l}T)$
$R_{jk}$ is the Ricci tensor with trace $R$, $T_{kl}$ is the stress-energy tensor with trace $T$. The Bianchi identity $\nabla_{j}R^{j}{}_{k}=\frac{1}{2}\nabla_{k}R$ implies $\nabla_{j}T^{j}{}_{k}=0$. Solutions of the Einstein equations are solutions of the new theory.
Shortly after, we found a parametrization showing that Harada’s equations are equivalent to the Einstein equations modified by a supplemental conformal Killing tensor (CKT) that is also divergence-free [3]:
$R_{kl}-\tfrac{1}{2}Rg_{kl}=T_{kl}+K_{kl}$ (2)
$\nabla_{j}K_{kl}+\nabla_{k}K_{jl}+\nabla_{l}K_{jk}$ (3)
$=\tfrac{1}{6}(g_{kl}\nabla_{j}K+g_{jl}\nabla_{k}K+g_{jk}\nabla_{l}K)$
For this reason the theory was named Conformal Killing Gravity (CKG). The reformulation makes Harada’s extension of GR explicit through the conformal Killing term, that satisfies $\nabla^{k}K_{kl}=0$ and enters as a new source term in the equations.
Feng and Chen [4] proposed an action principle for CKG. Some references on geometrical and physical applications of conformal Killing tensors are [5, 6, 7, 8].
As $R_{kl}$ contains second order derivatives of the metric tensor, higher orders in the field equations (2) may arise with the tensor $K_{jk}$. This does not occur in the present work, as well as in [3], where we obtained a realization of the conformal Killing parametrization in FRW background. In [3] the CKT has the perfect fluid form, candidate for representing the dark sector, and contains the scale factor with no derivatives. The Friedmann equations are thus second order in the metric, and reproduced the same forecasting obtained by Harada [2] with eqs.(1). Vacuum cosmological solutions, wormhole and black hole solutions were obtained by Clément and Nouicer [9]. In FRW background, CKG is embraced by a general parametrization of Codazzi tensors [10].
In the next work [11] we deepened the geometrical aspects and the cosmological consequences. In particular we showed that the density contrast in the matter era behaved as in $\Lambda$CDM and provided a fit of the Hubble parameter versus redshift with CC and BAO, with forecast of future singularities.
Most papers on CKG dealt with a static background. Barnes [12] found the general spherically symmetric vacuum solution and [13, 14] the general solution with Maxwell source. Junior, Lobo and Rodriguez [15] investigated regular black holes solutions of CKG coupled to nonlinear electrodynamics and scalar fields. Further in [16] they explored black bounce solutions in CKG coupled to nonlinear electrodynamics and scalar fields. pp-wave solutions were studied by Barnes [17].
In this paper we show a realization of the conformal Killing parametrization in a static spherically symmetric background
$ds^{2}=-b^{2}(r)dt^{2}+f_{1}^{2}(r)dr^{2}+f_{2}^{2}(r)d\Omega_{2}^{2}$
There are clear advantages in the parametrization approach (2): the results based on the equations (1) are here reobtained with simplicity and covariantly.
In Section II we recollect the covariant description of the Ricci tensor of a spherically symmetric static space-time found in [18]. Then we show that an anisotropic conformal Killing tensor is naturally hosted in such spaces. It extends the perfect fluid form recently shown by Barnes in [17] in spherical symmetry.
In Section III we write the field equations of CKG for an anisotropic fluid source: they are second order differential equations in the metric functions, as in GR.
An interesting result for the metric coefficients is obtained: if $f_{2}=r$ then $p_{r}+\mu=0$ if and only if $bf_{1}=(\kappa_{3}r^{2}+\kappa_{4})^{-1/2}$. It extends the result obtained by Barnes [13] and Clément and Noucier [9] for vacuum spacetimes and linear electrodynamics.
In section IV the vacuum case is analyzed. We prove the equivalence of the third order differential equation (26) for $b^{2}(r)$ by Barnes with the second order one obtained by us, eq.(25). The first one descends from Harada’s equations (1), the latter descends from (2).
We show that the (unique) Schwarzschild-like solution is the vacuum by Harada and remark that it cannot originate from a perfect fluid CKT.
In section V we write the CKG equations for nonlinear electrodynamics. The same result for the metric functions in Sect. III is proven in the nonlinear case.
In the linear case (electric and magnetic monopoles) the general solution by Barnes [13] and Clément and Noucier [9] is here obtained as solution of a second order equation.
For the purely magnetic case, we write the second order equation (44) for $b^{2}$ with generic magnetic Lagrangian. It straightforwardly reproduces the Lagrangians obtained for black-hole solutions by Junior et al. [15], and a generalization of the Hayward metric.
II Anisotropic CKT
A covariant characterization [20, 19, 18] of static spacetimes is the existence of a time-like unit vector field that is shear, expansion, vorticity free, with closed acceleration $\dot{u}_{j}=u^{k}\nabla_{k}u_{j}$:
$\nabla_{k}u_{j}=-u_{k}\dot{u}_{j},\qquad\nabla_{j}\dot{u}_{k}=\nabla_{k}\dot{u}_{j}$ (4)
With $\eta=\dot{u}^{k}\dot{u}_{k}$ the normalized acceleration vector is
$\chi_{k}=\frac{\dot{u}_{k}}{\sqrt{\eta}}$ (5)
We focus on static spherically-symmetric spacetimes
$ds^{2}=-b^{2}(r)dt^{2}+f_{1}^{2}(r)dr^{2}+f_{2}^{2}(r)(d\theta^{2}+\sin^{2}\theta d\phi^{2})$ (6)
In this (comoving) frame: $u_{0}=-b$, $u_{\mu}=0$, $\dot{u}_{0}=0$, $\dot{u}_{r}=b^{\prime}/b$, $\dot{u}_{\theta}=\dot{u}_{\phi}=0$. A prime denotes a derivative in the variable $r$. Then:
$\sqrt{\eta}=\frac{b^{\prime}}{f_{1}b}$ (7)
and $\chi_{0}=0$, $\chi_{r}=f_{1}$, $\chi_{\theta}=\chi_{\phi}=0$
The covariant expression of the Ricci tensor was obtained in [18], eq.85:
$R_{kl}=\frac{R+4\nabla_{p}\dot{u}^{p}}{3}u_{k}u_{l}+\frac{R+\nabla_{p}\dot{u}^{p}}{3}g_{kl}$ (8)
$+\Sigma(r)\left[\chi_{k}\chi_{l}-\frac{u_{k}u_{l}+g_{kl}}{3}\right]$
where the coefficients are (eqs. 87, 89 of [18]):
$\nabla_{p}\dot{u}^{p}=$ $\frac{1}{bf_{1}^{2}}\left[b^{\prime\prime}-b^{\prime}\left(\frac{f_{1}^{\prime}}{f_{1}}-2\frac{f_{2}^{\prime}}{f_{2}}\right)\right],$ (9)
$\Sigma(r)=$ $-\frac{1}{bf_{1}^{2}}\left[b^{\prime\prime}-b^{\prime}\left(\frac{f_{1}^{\prime}}{f_{1}}+\frac{f_{2}^{\prime}}{f_{2}}\right)\right]$ (10)
$-\frac{1}{f_{1}^{2}}\left[\frac{f_{1}^{2}}{f_{2}^{2}}+\frac{f_{2}^{\prime\prime}}{f_{2}}-\left(\frac{f_{2}^{\prime}}{f_{2}}\right)^{2}-\frac{f_{1}^{\prime}f_{2}^{\prime}}{f_{1}f_{2}}\right],$
$R=R^{k}{}_{k}$ and the curvature $R^{\star}$ of the space-like submanifold are (eqs. 90, 91 of [18]):
$R=R^{\star}-2\nabla_{p}\dot{u}^{p},$ (11)
$R^{\star}=\frac{2}{f_{2}^{2}}-\frac{2}{f_{1}^{2}}\left[2\frac{f_{2}^{\prime\prime}}{f_{2}}-2\frac{f_{1}^{\prime}}{f_{1}}\frac{f_{2}^{\prime}}{f_{2}}+\left(\frac{f_{2}^{\prime}}{f_{2}}\right)^{2}\right].$ (12)
In the background (6) we consider a symmetric tensor with the anisotropic structure of the Ricci tensor:
$K_{kl}=\mathsf{A}(r)u_{k}u_{l}+\mathsf{B}(r)g_{kl}+\mathsf{C}(r)\chi_{k}\chi_{l}$ (13)
The following result shows that a static spherically-symmetric spacetime always hosts an anisotropic conformal Killing tensor (the proof is in Appendix 2):
Theorem 1
In the static spherically symmetric spacetime (6) the tensor (13) is a divergence-free conformal Killing tensor if and only if
$\mathsf{A}=\kappa_{2}f_{2}^{2}-2\kappa_{3}b^{2},$
$\mathsf{B}=\kappa_{1}+2\kappa_{2}f_{2}^{2}+\kappa_{3}b^{2}$
$\mathsf{C}=-\kappa_{2}f_{2}^{2}$
where $\kappa_{1}$, $\kappa_{2}$ and $\kappa_{3}$ are arbitrary constants.
$\mathsf{A}$, $\mathsf{B}$ and $\mathsf{C}$ are combinations of the metric functions $b$, $f_{1}$, $f_{2}$ with no derivatives. Therefore, the CKG equations (2) and (3) contain derivatives of the metric not higher than second order.
Example 2 (Perfect fluid CKT)
$Z_{j}=bu_{j}$ is a Killing vector, $\nabla_{j}Z_{k}+\nabla_{k}Z_{j}=0.$ Following the strategy proposed in [7], $K_{jk}=\alpha Z_{j}Z_{k}+\beta\,g_{jk}$ with constant $\alpha$ and scalar function $\beta$, is a CKT. The divergence-free condition is $\beta=-\frac{1}{2}\alpha b^{2}+\gamma$.
$K_{jk}=\alpha b^{2}(u_{j}u_{k}-\tfrac{1}{2}g_{jk})+\gamma g_{jk}$
is the divergence-free CKT recently shown by Barnes in [17]. It is a particular case of Theorem 1 with $\kappa_{2}=0$.
III CKG for the anisotropic fluid
The structures of the Ricci tensor (8) and of the CKT (13) are compatible with the stress-energy tensor of an anisotropic fluid without heat flow:
$T_{kl}=(\mu+p_{\perp})u_{k}u_{l}+p_{\perp}g_{kl}+(p_{r}-p_{\perp})\chi_{k}\chi_{l}$ (14)
$\mu$ is the energy density, $p_{r}$ and $p_{\perp}$ are the radial and transversal pressures in the comoving frame, $\chi_{k}=\dot{u}_{k}/\sqrt{\eta}$. The tensor is rewritten as
$T_{kl}=(\mu+P)u_{k}u_{l}+Pg_{kl}+(p_{r}-p_{\perp})\left[\chi_{k}\chi_{l}-\frac{g_{kl}+u_{k}u_{l}}{3}\right]$
where $P=\frac{1}{3}(p_{r}+2p_{\perp})$ is the total pressure. The CKG equations with the fluid source are
$\frac{R+4\nabla_{p}\dot{u}^{p}}{3}u_{k}u_{l}+\frac{2\nabla_{p}\dot{u}^{p}-R}{6}g_{kl}+\Sigma\left[\chi_{k}\chi_{l}-\dfrac{u_{k}u_{l}+g_{kl}}{3}\right]$
$=(\mu+P)u_{k}u_{l}+Pg_{kl}+(p_{r}-p_{\perp})\left[\chi_{k}\chi_{l}-\dfrac{u_{k}u_{l}+g_{kl}}{3}\right]$
$+\left[\mathsf{A}+\dfrac{\mathsf{C}}{3}\right]u_{k}u_{l}+\left[\mathsf{B}+\dfrac{\mathsf{C}}{3}\right]g_{kl}+\mathsf{C}\left[\chi_{k}\chi_{l}-\dfrac{u_{k}u_{l}+g_{kl}}{3}\right]$
They give three scalar equations:
$\frac{1}{3}(R+4\nabla_{p}\dot{u}^{p})=(\mu+P)+\mathsf{A}+\frac{1}{3}\mathsf{C}$
$\frac{1}{6}(2\nabla_{p}\dot{u}^{p}-R)=P+\mathsf{B}+\frac{1}{3}\mathsf{C}$
$\Sigma=(p_{r}-p_{\perp})+\mathsf{C}$
Rearranging terms and using $R=R^{\star}-2\nabla_{p}\dot{u}^{p}$ and Theorem 1 we obtain
Proposition 3
The field equations of CKG in the static spherically symmetric metric (6) with the anisotropic CKT eq.(13) are:
$\frac{R^{\star}}{2}$ $=\mu-3\kappa_{3}b^{2}-\kappa_{2}f_{2}^{2}-\kappa_{1}$
$\nabla_{p}\dot{u}^{p}$ $=\frac{3}{2}P+\frac{1}{2}\mu+2\kappa_{2}f_{2}^{2}+\kappa_{1}$ (15)
$\Sigma$ $=(p_{r}-p_{\perp})+\mathsf{C}$
Remark 4
Let $f_{2}=r$ and consider eqs.(15) for a perfect fluid ($p_{r}=p_{\perp}$). In GR they are $\mu=\frac{1}{2}R^{\star}$, $P=\frac{2}{3}\nabla_{p}\dot{u}^{p}-\frac{1}{6}R^{\star}$ and $\Sigma=0$: three equations for the unknowns $\mu$, $p$, $b$ and $f_{1}$. A further condition, such as an equation of state, is needed. The same occurs for the conformal Killing equations.
III.1 Properties of metric functions
Hereafter we consider the static spherical metric with $f_{2}=r$, and set $f_{1}=h/b$.
$ds^{2}=-b^{2}(r)dt^{2}+\frac{h^{2}(r)}{b^{2}(r)}dr^{2}+r^{2}d\Omega^{2}$ (16)
We prove a remarkable property of the metric function $h(r)$. First we assert the following geometric result.
Lemma 5
$\frac{R^{\star}}{2}+\nabla_{p}\dot{u}^{p}+\Sigma=\dfrac{3b^{2}}{r}\dfrac{h^{\prime}}{h^{3}}$ (17)
Proof. Using (9), (10) and (12) the following relations are straigthforwardly obtained:
$R^{\star}=\dfrac{2}{r^{2}}+\dfrac{4b^{2}}{r}\dfrac{h^{\prime}}{h^{3}}-\dfrac{4bb^{\prime}}{rh^{2}}-\dfrac{2}{r^{2}}\dfrac{b^{2}}{h^{2}}$ (18)
$\nabla_{p}\dot{u}^{p}=\dfrac{1}{h^{2}}(b^{\prime 2}+bb^{\prime\prime})-\dfrac{bb^{\prime}}{h^{2}}\dfrac{h^{\prime}}{h}+\dfrac{2bb^{\prime}}{rh^{2}}$ (19)
$\Sigma=-\dfrac{1}{h^{2}}(b^{\prime 2}+bb^{\prime\prime})+bb^{\prime}\dfrac{h^{\prime}}{h^{3}}+\dfrac{1}{r^{2}}\dfrac{b^{2}}{h^{2}}-\dfrac{1}{r^{2}}+\dfrac{b^{2}}{r}\dfrac{h^{\prime}}{h^{3}}$ (20)
The result (17) follows. $\square$
We are ready to prove
Proposition 6
Consider CKG in the metric (16) with anisotropic fluid source. Then:
$\frac{1}{h^{2}}=\kappa_{3}r^{2}+\kappa_{4}\quad\Longleftrightarrow\quad p_{r}=-\mu$ (21)
Proof. Addition of eqs. (15) with $\mathsf{C}=-\kappa_{2}r^{2}$ and $P=\frac{1}{3}(p_{r}+2p_{\perp})$ gives
$\frac{R^{\star}}{2}+\nabla_{p}\dot{u}^{p}+\Sigma=\frac{3}{2}(p_{r}+\mu)-3\kappa_{3}b^{2}$
With (17) we obtain
$\frac{2b^{2}}{r}\frac{h^{\prime}}{h^{3}}+2\kappa_{3}b^{2}=p_{r}+\mu$
If $p_{r}+\mu=0$ it is $\dfrac{h^{\prime}}{h^{3}}=-\kappa_{3}r$ that integrates to (21) with $\kappa_{4}$ a constant. Conversely if $h^{-2}=\kappa_{3}r^{2}+\kappa_{4}$ then $p_{r}+\mu=0$. $\square$
If $p_{r}+\mu=0$ it is
$f_{1}(r)=\dfrac{1}{b(r)\sqrt{\kappa_{3}r^{2}+\kappa_{4}}}$ (22)
Eq.(22) was obtained by Barnes [12] while investigating vacuum solutions and also in solving CKG field equations of linear electrodynamics [13].
IV Vacuum solutions
In absence of matter eqs.(15) become
$R^{\star}$ $=-6\kappa_{3}b^{2}-2\kappa_{2}r^{2}-2\kappa_{1}$
$\nabla_{p}\dot{u}^{p}$ $=2\kappa_{2}r^{2}+\kappa_{1}$ (23)
$\Sigma$ $=\mathsf{C}=-\kappa_{2}r^{2}$
It is also $h^{-2}=\kappa_{3}r^{2}+\kappa_{4}$. The remaining metric function $b^{2}$ is determined by $\Sigma=\mathsf{C}$. With the substitution $y=b^{2}$ the anisotropic term (20) now is
$\Sigma=-(\kappa_{3}r^{2}+\kappa_{4})\frac{y^{\prime\prime}}{2}-\kappa_{3}r\frac{y^{\prime}}{2}+\kappa_{4}\frac{y}{r^{2}}-\frac{1}{r^{2}}$ (24)
and the equation $\Sigma=-\kappa_{2}r^{2}$ for $y(r)$ is:
$(\kappa_{3}r^{2}+\kappa_{4})y^{\prime\prime}+\kappa_{3}ry^{\prime}-2\frac{\kappa_{4}}{r^{2}}y+\frac{2}{r^{2}}=2\kappa_{2}r^{2}$ (25)
Proposition 7
The second order equation (25) is equivalent to the third-order equation obtained by Barnes in [12] eq.(16), with computer algebra:
$(\kappa_{3}r^{2}+\kappa_{4})$ $r^{3}y^{\prime\prime\prime}+(\kappa_{3}r^{2}-2\kappa_{4})r^{2}y^{\prime\prime}$
$-(\kappa_{3}r^{2}+2\kappa_{4})ry^{\prime}+8\kappa_{4}y=8.$ (26)
Proof. The equation $\Sigma=\mathsf{C}$ gives $\Sigma^{\prime}=\mathsf{C}^{\prime}$. Being $\mathsf{C}=-\kappa_{2}r^{2}$ it is $\mathsf{C}^{\prime}=\frac{2}{r}\mathsf{C}$. Thus the anisotropic term satisfies $\Sigma^{\prime}=\frac{2}{r}\Sigma$, and the integral is $\Sigma=\mathsf{C}$:
$\Sigma^{\prime}-\dfrac{2}{r}\Sigma=0\quad\Longleftrightarrow\quad\Sigma=\mathsf{C}$ (27)
The left hand side of (27) is evaluated
$\Sigma^{\prime}-\dfrac{2}{r}\Sigma=-\dfrac{y^{\prime\prime\prime}}{2}(\kappa_{3}r^{2}+\kappa_{4})-y^{\prime\prime}\left(\dfrac{\kappa_{3}}{2}r-\dfrac{\kappa_{4}}{r}\right)$
$+y^{\prime}\left(\dfrac{\kappa_{3}}{2}+\dfrac{\kappa_{4}}{r^{2}}\right)-y\dfrac{4\kappa_{4}}{r^{3}}+\dfrac{4}{r^{3}}$
and entails the equation by Barnes. $\square$
The second order equation (25) is solved ($x=r\sqrt{\kappa}_{3}$, $\kappa_{4}=1$): the sum of the homogeneous (with coefficients $c_{1,2}$) and the inhomogeneous solutions
$b^{2}(r)=$ $c_{1}\frac{\sqrt{1+x^{2}}}{x}+c_{2}\left[\frac{\sqrt{1+x^{2}}}{x}{\rm arcsh}{x}-1\right]$ (28)
$+1-\frac{3\kappa_{2}}{2\kappa_{3}^{2}}\left[\frac{\sqrt{1+x^{2}}}{x}{\rm arcsh}{x}-1-\frac{x^{2}}{3}\right]$
Example 8 (Case $\mathbf{h=1}$)
The vacuum static spherical solution of CKG with $h=1$ (i.e. $\kappa_{3}=0$ and $\kappa_{4}=1$) was found by Harada [1], and extends the Schwarzschild-de Sitter solution of GR:
$b^{2}=1-\dfrac{2M}{r}-\dfrac{\Lambda}{3}r^{2}-\dfrac{\lambda}{5}r^{4}$ (29)
The following result holds:
Proposition 9
Eq.(29) is the unique solution with $h=1$ of the vacuum equations (23) with $\kappa_{1}=-\Lambda$, $\kappa_{2}=-\lambda$.
Proof. With $\kappa_{3}=0$ and $\kappa_{4}=1$ in (25), the equation $\Sigma=\mathsf{C}$ is the Euler equation $r^{2}y^{\prime\prime}-2y=-2+2\kappa_{2}r^{4}$ with solution
$y=b^{2}=1+\dfrac{c_{1}}{r}+c_{2}r^{2}+\dfrac{1}{5}\kappa_{2}r^{4}$ (30)
Thus $c_{1}=-2M$, $c_{2}=-\frac{1}{3}\Lambda$ and $\lambda=-\kappa_{2}$.
The solution must also solve the other equations in (23): $R^{\star}=-2\kappa_{2}r^{2}-2\kappa_{1}$ and $\nabla_{p}\dot{u}^{p}=2\kappa_{2}r^{2}+\kappa_{1}$.
With $h=1$ and $b^{2}=y$ eqs.(18) and (19) become:
$R^{\star}=-\frac{2}{r}y^{\prime}-\frac{2}{r^{2}}y+\frac{2}{r^{2}}=2\lambda r^{2}+2\Lambda$
$\nabla_{p}\dot{u}^{p}=\frac{1}{2}y^{\prime\prime}+\frac{1}{r}y^{\prime}=-2\lambda r^{2}-\Lambda$
The equations are satisfied with $\kappa_{1}=-\Lambda$. $\square$
The Schwarzschild de-Sitter metric occurs if and only if $\kappa_{2}=0$, i.e. $\mathsf{C}=0$, giving the perfect fluid CKT. Then, a perfect fluid CKT in the field equations cannot originate Harada’s vacuum solution (29).
V (non)Linear electrodynamics
in CKG
In [18] we specified the covariant form of the stress-energy tensor of nonlinear electrodynamics in static spherically-symmetric spacetimes:
$T_{jk}^{nlin}=2(\mathbb{E}^{2}+\mathbb{B}^{2})\left[u_{j}u_{k}-\chi_{j}\chi_{k}\right]\mathscr{L}_{F}(F)$ (31)
$+2g_{jk}[\mathbb{B}^{2}\mathscr{L}_{F}(F)-\mathscr{L}(F)].$
$F=\tfrac{1}{4}F_{jk}F^{jk}$ is the Faraday scalar, $\mathscr{L}_{F}=d\mathscr{L}/dF.$ In the static setting it is $F=\tfrac{1}{2}(\mathbb{B}^{2}-\mathbb{E}^{2})$, where
Proposition 10
$\mathbb{E}(r)=\dfrac{q_{e}}{r^{2}\mathscr{L}_{F}(F)},\quad\mathbb{B}=\frac{q_{m}}{r^{2}}$ (32)
Proof. Eq.28 in [18] is: $\nabla_{j}[\chi^{j}\mathbb{E}\mathscr{L}_{F}]=\sqrt{\eta}\mathbb{E}\mathscr{L}_{F}$. Explicitly, using (7) and the formulas in Appendix 1
$\left(\frac{2}{rf_{1}}+\frac{b^{\prime}}{f_{1}b}\right)(\mathbb{E}\mathscr{L}_{F})+\frac{1}{f_{1}}\frac{d}{dr}(\mathbb{E}\mathscr{L}_{F})=\frac{b^{\prime}}{f_{1}b}(\mathbb{E}\mathscr{L}_{F})$
Terms cancel and the linear equation gives the first result, where $q_{e}$ is an integration constant.
Eq.18 in [18] is $\sqrt{\eta}\chi^{p}\nabla_{p}\mathbb{B}=\mathbb{B}\left[\nabla_{p}\dot{u}^{p}-\eta-\frac{\chi^{s}\nabla_{s}\eta}{2\sqrt{\eta}}\right]$. Using (56) the equation is integrated and yields the monopole solution. $\square$
(different deductions of the proposition are in [21, 22]).
The tensor structure of the Ricci and of the stress-energy tensors is matched by the conformal Killing tensor. Thus the field equations of CKG are:
$\frac{R+4\nabla_{p}\dot{u}^{p}}{3}u_{k}u_{l}+\frac{2\nabla_{p}\dot{u}^{p}-R}{6}g_{kl}$
$+\Sigma(r)\left[\chi_{k}\chi_{l}-\frac{u_{k}u_{l}+g_{kl}}{3}\right]=\frac{4}{3}(\mathbb{E}^{2}+\mathbb{B}^{2})\mathscr{L}_{F}(F)u_{k}u_{l}$
$+2\left[\dfrac{1}{3}(2\mathbb{B}^{2}-\mathbb{E}^{2})\mathscr{L}_{F}(F)-\mathscr{L}(F)\right]g_{kl}$
$-2(\mathbb{E}^{2}+\mathbb{B}^{2})\mathscr{L}_{F}(F)\left[\chi_{k}\chi_{l}-\frac{u_{k}u_{l}+g_{kl}}{3}\right]$
$+\left(\mathsf{A}+\dfrac{\mathsf{C}}{3}\right)u_{k}u_{l}+\left(\mathsf{B}+\dfrac{\mathsf{C}}{3}\right)g_{kl}+\mathsf{C}\left[\chi_{k}\chi_{l}-\frac{u_{k}u_{l}+g_{kl}}{3}\right]$
By equating the coefficients one obtains three scalar equations. A rearrangement of terms and use of
$R=R^{\star}-2\nabla_{p}\dot{u}^{p}$ give
$\frac{1}{2}R^{\star}=2\mathscr{L}_{F}(F)\mathbb{E}^{2}+2\mathscr{L}(F)+\mathsf{A}-\mathsf{B}$
$\nabla_{p}\dot{u}^{p}=2\mathscr{L}_{F}(F)\mathbb{B}^{2}-2\mathscr{L}(F)+\frac{\mathsf{A}}{2}+\mathsf{B}+\frac{\mathsf{C}}{2}$ (33)
$\Sigma=-2(\mathbb{E}^{2}+\mathbb{B}^{2})\mathscr{L}_{F}(F)+\mathsf{C}$
Also in this case we are able to show the validity of (22). The following result holds:
Proposition 11
Consider CKG coupled with nonlinear electrodynamics in the metric (16). Then
$\frac{1}{h^{2}}=\kappa_{3}r^{2}+\kappa_{4}$
Proof. The sum of eqs. (33) is $\tfrac{1}{2}R^{\star}+\nabla_{p}\dot{u}^{p}+\Sigma=\tfrac{3}{2}(\mathsf{A}+\mathsf{C})$. Now use Lemma 5 and $\mathsf{A}+\mathsf{C}=-2\kappa_{3}b^{2}$. $\square$
V.1 Linear electrodynamics
In linear electrodynamics $\mathscr{L}(F)=F=\tfrac{1}{2}(\mathbb{B}^{2}-\mathbb{E}^{2})$. The electric field is Coulomb $\mathbb{E}(r)=q_{e}/r^{2}$, so that
$\mathbb{E}^{2}+\mathbb{B}^{2}=\frac{q^{2}}{r^{4}}$
with $q^{2}=q_{e}^{2}+q_{m}^{2}$. Equations (33) take the form
$\frac{1}{2}R^{\star}=\frac{q^{2}}{r^{4}}-3\kappa_{3}b^{2}-\kappa_{2}r^{2}-\kappa_{1}$ (34)
$\nabla_{p}\dot{u}^{p}=\frac{q^{2}}{r^{4}}+2\kappa_{2}r^{2}+\kappa_{1}$ (35)
$\Sigma=-2\frac{q^{2}}{r^{4}}+\mathsf{C}$ (36)
$\Sigma$ is given by (24) and $\mathsf{C}=-\kappa_{2}r^{2}$. Eq.(36) is the second order differential equation for $y=b^{2}$:
$(\kappa_{3}r^{2}+\kappa_{4})y^{\prime\prime}+\kappa_{3}ry^{\prime}-2\kappa_{4}\dfrac{y}{r^{2}}+\dfrac{2}{r^{2}}=2\kappa_{2}r^{2}+\dfrac{4q^{2}}{r^{4}}$ (37)
Proposition 12
In CKG coupled with linear electrodynamics the second order equation for $b^{2}(r)$ in a static spherically symmetric background is equivalent to the following third order equation:
$(\kappa_{3}r^{2}+\kappa_{4})r^{3}y^{\prime\prime\prime}+(\kappa_{3}r^{2}-2\kappa_{4})r^{2}y^{\prime\prime}$ (38)
$\qquad-(\kappa_{3}r^{2}+2\kappa_{4})ry^{\prime}+8\kappa_{4}y=8-\dfrac{24q^{2}}{r^{2}}$
This is eq.14 worked by Barnes in [13] and the “master equation” by Clément and Noucier [9].
Proof. As in the vacuum case, $\Sigma=-2\frac{q^{2}}{r^{4}}+\mathsf{C}$ gives $\Sigma^{\prime}=\frac{8q^{2}}{r^{5}}+\mathsf{C}^{\prime}$. Since $\mathsf{C}=-\kappa_{2}r^{2}$ it is $\mathsf{C}^{\prime}=\frac{2}{r}\mathsf{C}=\frac{2}{r}\left[\Sigma+2\frac{q^{2}}{r^{4}}\right]$. Thus the anisotropic term satisfies $\Sigma^{\prime}-\frac{2}{r}\Sigma=\frac{12q^{2}}{r^{5}}$; on the other hand the integral is $\Sigma=-2\frac{q^{2}}{r^{4}}+\mathsf{C}$. The following equivalence holds:
$\Sigma^{\prime}-\dfrac{2}{r}\Sigma=\dfrac{12q^{2}}{r^{5}}\quad\Longleftrightarrow\quad\Sigma=-2\dfrac{q^{2}}{r^{4}}+\mathsf{C}$ (39)
Using (24) the left hand side of (39) is evaluated and gives the third order equation. $\square$
As a check, the second order equation (37) is solved for $\kappa_{4}=1$, $x=r\sqrt{\kappa}_{3}$. The solution, evaluated with Mathematica, coincides with that by Clément and Noucier [9]
$y(x)$ $=c_{1}\frac{\sqrt{1+x}}{x}+c_{2}\left[\frac{\sqrt{1+x^{2}}}{x}{\rm arcsh}(x)-1\right]+1$
$-\frac{3\kappa_{2}}{2\kappa_{3}^{2}}\left[\frac{\sqrt{1+x^{2}}}{x}{\rm arcsh}(x)-1-\frac{x^{2}}{3}\right]+\kappa_{3}q^{2}(2+\frac{1}{x^{2}})$
Example 13 (Case $\mathbf{h=1}$)
The metric function
$b^{2}(r)=1-\dfrac{2M}{r}-\dfrac{\Lambda}{3}r^{2}+\dfrac{q^{2}}{r^{2}}-\dfrac{\lambda}{5}r^{4}$ (40)
was obtained in [15, 13] with different strategies. In [15] Junior, Lobo and Rodriguez posed a functional form of $b(r)$ including the term $-\frac{\lambda}{5}r^{4}$ that characterizes CKG, and then engineered a nonlinear electrodynamics Lagrangian and its derivative. On the other hand, in [13] Barnes used computer algebra to obtain the same and more general solutions.
Here the solution (40) is achieved in a much simpler way:
Proposition 14
The unique solution of (34)–(36) with $h=1$ (i.e. $\kappa_{3}=0$, $\kappa_{4}=1$) is the metric function (40), with $\kappa_{1}=-\Lambda$, $\kappa_{2}=-\lambda$.
Proof. The equation $\Sigma=-2\frac{q^{2}}{r^{4}}+\mathsf{C}$ with $\Sigma$ in (24) and $y=b^{2}$ is Euler’s equation $r^{2}y^{\prime\prime}-2y+2=4\frac{q^{2}}{r^{2}}+2\kappa_{2}r^{4}$ with solution
$y=b^{2}=1+\frac{c_{1}}{r}+c_{2}r^{2}+\frac{q^{2}}{r^{2}}+\dfrac{\kappa_{2}}{5}r^{4}.$ (41)
This is (40) with $c_{1}=-2M$, $c_{2}=-\frac{\Lambda}{3}$ and $\lambda=-\kappa_{2}$.
Let’s show that it also solves (34) and (35):
$\tfrac{1}{2}R^{\star}=q^{2}/r^{4}-\kappa_{2}r^{2}-\kappa_{1}$
$\nabla_{p}\dot{u}^{p}=q^{2}/r^{4}+2\kappa_{2}r^{2}+\kappa_{1}$
With (40), $h=1$ and $b^{2}=y$, eqs.(18) and (19) give:
$\tfrac{1}{2}R^{\star}=\tfrac{1}{r^{2}}-\tfrac{1}{r}y^{\prime}-\tfrac{1}{r^{2}}y=q^{2}/r^{4}-\kappa_{2}r^{2}-3c_{2}$
$\nabla_{p}\dot{u}^{p}=\tfrac{1}{2}y^{\prime\prime}+\tfrac{1}{r}y^{\prime}=q^{2}/r^{4}+2\kappa_{2}r^{2}+3c_{2}$
Equality is achieved with $\kappa_{1}=3c_{2}$. $\square$
Example 15
For $\kappa_{3}=1$ and $\kappa_{4}=0$ eq.(37) is solved by
$y=\lambda-\dfrac{\Lambda}{3}r^{2}+m\ln r-\dfrac{1}{2r^{2}}+\dfrac{q^{2}}{4r^{4}}$ (42)
This is eq. 19 in Barnes or 2.21 in Clément and Nouicer.
V.2 Nonlinear electrodynamics:
purely magnetic solutions
In this case $\mathbb{E}(r)=0$ and $\mathbb{B}=q_{m}/r^{2}$ (magnetic monopole). Then $F=\frac{\mathbb{B}^{2}}{2}=\frac{q_{m}^{2}}{2r^{4}}.$ The third equation of (33) is
$\Sigma=-2\dfrac{q_{m}^{2}}{r^{4}}\mathscr{L}_{F}(F)+\mathsf{C}$
Now $\mathscr{L}_{F}=\dfrac{d\mathscr{L}}{dF}=\dfrac{d\mathscr{L}}{dr}\dfrac{dr}{dF}=-\mathscr{L}^{\prime}\frac{r^{5}}{2q_{m}^{2}}$, so that the previous equation rewrites as
$\Sigma=r\mathscr{L}^{\prime}+\mathsf{C}.$ (43)
The knowledge of $\mathscr{L}(r)$ in (43) gives a second order differential equation for the metric function $b^{2}=y$:
$(\kappa_{3}r^{2}+\kappa_{4})y^{\prime\prime}+\kappa_{3}ry^{\prime}-\frac{2\kappa_{4}}{r^{2}}y+\frac{2}{r^{2}}=-2(r\mathscr{L}^{\prime}-\kappa_{2}r^{2})$ (44)
Example 16 (Case $\mathbf{h=1}$)
For $\kappa_{3}=0$ and $\kappa=1$ the solution by the method of variation of parameters is:
[…]
Given ${\mathscr{L}}(F)$, $F=q_{m}^{2}/(2r^{4})$, one evaluates $b(r)$.
In [15] a different strategy is pursued: an input metric with faithful properties (such as regularity in the origin) is chosen, and the corresponding Lagrangian is reconstructed.
Here this procedure is greatly facilitated by eq.(43). It is illustrated to reproduce two interesting examples:
$\bullet$ Metric eq.34 in [15]:
$y=1-\frac{2M}{r}-\frac{\Lambda}{3}r^{2}+\frac{q_{m}^{2}}{r^{2}}-\frac{\lambda}{5}r^{4}+\frac{k_{0}}{r^{4}}$ (46)
With $\Sigma=\dfrac{y-1}{r^{2}}-\dfrac{1}{2}y^{\prime\prime}$ we get $\Sigma=\lambda r^{2}-\dfrac{2q_{m}^{2}}{r^{4}}-9\dfrac{k_{0}}{r^{6}}$. Eq.(43) is easily solved by
$\mathscr{L}$ $=\phi_{0}+\frac{q_{m}^{2}}{2r^{4}}+\phi_{1}r^{2}+3\frac{k_{0}}{2r^{6}}$
$=\phi_{0}+F+\phi_{1}\frac{q_{m}}{\sqrt{2F}}+3\frac{k_{0}(2F)^{3/2}}{2q_{m}^{3}}$ (47)
with constants $\phi_{0}$, $\phi_{1}$. This is eq 41 in [15].
$\bullet$ Bardeen-type metric eq.72 in [15]):
$y=1-\frac{2Mr^{2}}{(q_{m}^{2}+r^{2})^{3/2}}-\frac{\Lambda}{3}r^{2}-\frac{\lambda}{5}r^{4}$ (48)
Now $\Sigma=\lambda r^{2}-\dfrac{15Mr^{2}}{(q_{m}^{2}+r^{2})^{7/2}}$. Eq.(43) is easily solved:
$\mathscr{L}=\dfrac{3Mq_{m}^{2}}{(q_{m}^{2}+r^{2})^{5/2}}+\phi_{1}r^{2}+\phi_{0}$ (49)
with $\phi_{0}$ and $\phi_{1}$ constants. This is eq.73 in [15]. With $F=q_{m}^{2}/(2r^{4})$, $\mathscr{L}$ can be rewritten as
$\mathscr{L}(F)=\frac{3Mq_{m}^{2}}{\left(q_{m}^{2}+\dfrac{q_{m}}{\sqrt{2F}}\right)^{5/2}}+\phi_{1}\dfrac{q_{m}}{\sqrt{2F}}+\phi_{0}$ (50)
This is eq.77 presented in [15].
This is a new example:
$\bullet$ Hayward-like solution. The metric
$b^{2}=1-\frac{2Mr^{2}}{q_{m}^{3}+r^{3}}-\frac{\Lambda}{3}r^{2}-\frac{\lambda}{5}r^{4}$ (51)
is the Hayward black-hole when $\lambda,\Lambda=0$ [23]. It is inferred that $\Sigma=\lambda r^{2}-\dfrac{18Mr^{3}q_{m}^{3}}{(q_{m}^{3}+r^{3})^{3}}.$ Then (43) is solved by
$\mathscr{L}=$ $\frac{3Mq_{m}^{3}}{(q_{m}^{3}+r^{3})^{2}}+\phi_{1}r^{2}+\phi_{0}$
$=$ $\frac{3M(2F)^{3/2}}{\left(1+(2Fq_{m}^{2})^{3/4}\right)^{2}}+\phi_{1}\frac{q_{m}}{\sqrt{2F}}+\phi_{0}$ (52)
VI Conclusions
We have shown that static spherically symmetric spacetimes naturally host an anisotropic divergence-free conformal Killing tensor (Theorem 1). This makes the parametrization (2) of the Harada equations as modified Einstein equations effective for such background.
The CKG equations can support an anisotropic fluid source as well as (non)-linear electrodynamics. In both cases the equations are second order. This is a great advantage with respect to the solution by components of the third order Harada equations found in the existing literature. We prove the equivalence of our second order equations with the third order ones. Our approach recovers several results obtained so far in a simple and covariant way, and gives some new ones.
Appendix 1. Useful formulae
We collect useful formulae for static spherical spacetimes, taken from [18].
A scalar function $F(r)$ has gradient in the radial direction, given by the unit spacelike vector $\chi$:
$\nabla_{j}F=\chi_{j}\chi^{k}\nabla_{k}F=\chi_{j}\frac{F^{\prime}}{f_{1}}$
(the prime is $d/dr$). In particular:
$\frac{\chi^{s}\nabla_{s}\eta}{2\sqrt{\eta}}=\frac{1}{2\sqrt{\eta}}\frac{1}{f_{1}}\frac{d}{dr}\left(\frac{b^{\prime 2}}{b^{2}f_{1}^{2}}\right)$
$=\frac{1}{f_{1}^{2}}\left[\frac{b^{\prime\prime}}{b}-\frac{b^{\prime 2}}{b^{2}}-\frac{b^{\prime}f_{1}^{\prime}}{bf_{1}}\right]$ (53)
The vector $\chi_{k}=\dot{u}_{k}/\sqrt{\eta}$ is normalized, $\chi^{k}\chi_{k}=1$.
$\nabla_{j}\chi_{k}=\frac{\nabla_{j}\dot{u}_{k}}{\sqrt{\eta}}-\frac{\chi^{s}\nabla_{s}\eta}{2\eta}\chi_{j}\chi_{k}$ (54)
Since $\dot{u}_{k}$ is closed, also $\chi_{j}$ is closed: $\nabla_{j}\chi_{k}=\nabla_{k}\chi_{j}$. A consequence is $\chi^{j}\nabla_{j}\chi_{k}=\chi^{j}\nabla_{k}\chi_{j}=0$.
The gradient of the acceleration is (eq.16 in [19]):
$\nabla_{j}\dot{u}_{k}=$ $-\eta u_{k}u_{j}+\frac{\chi^{s}\nabla_{s}\eta}{2\sqrt{\eta}}\chi_{j}\chi_{k}$ (55)
$+\frac{1}{2}N_{jk}\left[\nabla_{p}\dot{u}^{p}-\eta-\frac{\chi^{s}\nabla_{s}\eta}{2\sqrt{\eta}}\right]$
where $N_{jk}=g_{jk}+u_{j}u_{k}-\chi_{j}\chi_{k}$ is a projection.
Eq. 87 in [18], eqs.(7) and (53) give:
$\nabla_{p}\dot{u}^{p}-\eta-\frac{\chi^{s}\nabla_{s}\eta}{2\sqrt{\eta}}=\frac{2}{f_{1}^{2}}\left[\frac{b^{\prime}}{b}\frac{f_{2}^{\prime}}{f_{2}}\right]$ (56)
One then obtains:
$\nabla_{j}\chi_{k}=\frac{1}{f_{1}}\left[\frac{f_{2}^{\prime}}{f_{2}}-\frac{b^{\prime}}{b}\right]u_{k}u_{j}+\frac{f_{2}^{\prime}}{f_{1}f_{2}}(g_{jk}-\chi_{j}\chi_{k})$ (57)
$\nabla_{p}\chi^{p}=2\frac{f_{2}^{\prime}}{f_{1}f_{2}}+\frac{b^{\prime}}{f_{1}b}$ (58)
Appendix 2
Let $K_{kl}=\mathsf{A}u_{k}u_{l}+\mathsf{B}g_{kl}+\mathsf{C}\chi_{k}\chi_{l}$.
The Conformal Killing condition (3) with the static equation (4) for $u_{i}$ and the closedness of $\chi_{i}$ is:
$0=$ $\nabla_{i}(\mathsf{B}-\tfrac{1}{6}K)g_{jk}+(\nabla_{j}\mathsf{B}-\tfrac{1}{6}K)g_{ki}+(\nabla_{k}\mathsf{B}-\tfrac{1}{6}K)g_{ij}$
$+(\nabla_{i}\mathsf{A}-2\mathsf{A}\dot{u}_{i})u_{j}u_{k}+(\nabla_{j}\mathsf{A}-2\mathsf{A}\dot{u}_{j})u_{k}u_{i}$
$+(\nabla_{k}\mathsf{A}-2\mathsf{A}\dot{u}_{k})u_{i}u_{j}$
$+(\nabla_{i}\mathsf{C})\chi_{j}\chi_{k}+(\nabla_{j}\mathsf{C})\chi_{k}\chi_{i}+(\nabla_{k}\mathsf{C})\chi_{i}\chi_{j}$
$+2\mathsf{C}(\chi_{i}\nabla_{j}\chi_{k}+\chi_{j}\nabla_{k}\chi_{i}+\chi_{k}\nabla_{i}\chi_{j})$
Since $\mathsf{A}$ only depends on $r$, it is $\nabla_{i}\mathsf{A}=\chi_{i}\chi^{k}\nabla_{k}\mathsf{A}=\chi_{i}{\mathsf{A}}^{\prime}/f_{1}$, and similarly for $\mathsf{B}$ and $\mathsf{C}$ and $K$. The equation, multiplied by $f_{1}$ becomes:
$0=$ $(\mathsf{B}^{\prime}-\tfrac{1}{6}K^{\prime})(\chi_{i}g_{jk}+\chi_{j}g_{ki}+\chi_{k}g_{ij})$ (59)
$+$ $(\mathsf{A}^{\prime}-2\frac{b^{\prime}}{b}\mathsf{A})(\chi_{i}u_{j}u_{k}+\chi_{j}u_{k}u_{i}+\chi_{k}u_{i}u_{j})$
$+$ $3\mathsf{C}^{\prime}\chi_{i}\chi_{j}\chi_{k}+2f_{1}\mathsf{C}(\chi_{i}\nabla_{j}\chi_{k}+\chi_{j}\nabla_{k}\chi_{i}+\chi_{k}\nabla_{i}\chi_{j})$
Contraction with $\chi^{i}\chi^{j}\chi^{k}$:
$0=3\mathsf{B}^{\prime}+3{\mathsf{C}}^{\prime}-\tfrac{1}{2}K^{\prime}=\tfrac{1}{2}({\mathsf{A}}^{\prime}+2{\mathsf{B}}^{\prime}+5{\mathsf{C}}^{\prime})$ (60)
The identity is used to simplify $K^{\prime}$ and ${\mathsf{A}}^{\prime}$ from the equation:
$0=$ $(\mathsf{A}^{\prime}-2\frac{b^{\prime}}{b}\mathsf{A})(\chi_{i}u_{j}u_{k}+\chi_{j}u_{k}u_{i}+\chi_{k}u_{i}u_{j})$
$+$ $\mathsf{C}^{\prime}[3\chi_{i}\chi_{j}\chi_{k}-(\chi_{i}g_{jk}+\chi_{j}g_{ki}+\chi_{k}g_{ij})]$
$+$ $2f_{1}\mathsf{C}(\chi_{i}\nabla_{j}\chi_{k}+\chi_{j}\nabla_{k}\chi_{i}+\chi_{k}\nabla_{i}\chi_{j})$
Contraction with $u^{i}u^{j}\chi^{k}$: $0=(\mathsf{A}^{\prime}-2\frac{b^{\prime}}{b}\mathsf{A})+\mathsf{C}^{\prime}+2f_{1}\mathsf{C}u^{j}\dot{\chi}_{j}$. It is $u^{j}\dot{\chi}_{j}=-\dot{u}^{j}\chi_{j}=-\sqrt{\eta}=-b^{\prime}/(bf_{1})$. Then: $\mathsf{A}^{\prime}+\mathsf{C}^{\prime}=2\frac{b^{\prime}}{b}(\mathsf{A}+\mathsf{C})$, with solution $\mathsf{A}+\mathsf{C}=-2\kappa_{3}b^{2}(r)$.
Contraction with $g^{ij}\chi^{k}$:
$0=$ $-\mathsf{A}^{\prime}+2\frac{b^{\prime}}{b}\mathsf{A}-3\mathsf{C}^{\prime}+2f_{1}\mathsf{C}\nabla^{j}\chi_{j}$
Use (58) and obtain:
$0=$ $-(\mathsf{A}^{\prime}+\mathsf{C}^{\prime})+2\frac{b^{\prime}}{b}(\mathsf{A}+\mathsf{C})-2\mathsf{C}^{\prime}+4\frac{f_{2}^{\prime}}{f_{2}}\mathsf{C}$
It follows that ${\mathsf{C}}=-\kappa_{2}f^{2}_{2}$. Then ${\mathsf{A}}=-2\kappa_{3}b^{2}+\kappa_{2}f^{2}_{2}$. Eq.(60) gives ${\mathsf{B}}=\kappa_{3}b^{2}+2\kappa_{2}f^{2}_{2}+\kappa_{1}$ and $K=6\kappa_{3}b^{2}+6\kappa_{2}f^{2}_{2}+4\kappa_{1}$.
The found parameters $\mathsf{A,B,C}$ are inserted in (59). Up to a factor $2\kappa_{2}f_{1}f_{2}^{2}$ it is:
$0=$ $\frac{f_{2}^{\prime}}{f_{1}f_{2}}(\chi_{i}g_{jk}+\chi_{j}g_{ki}+\chi_{k}g_{ij})$ (61)
$+$ $\left[\frac{f_{2}^{\prime}}{f_{1}f_{2}}-\frac{b^{\prime}}{bf_{1}}\right](\chi_{i}u_{j}u_{k}+\chi_{j}u_{k}u_{i}+\chi_{k}u_{i}u_{j})$
-- $3\frac{f_{2}^{\prime}}{f_{2}f_{1}}\chi_{i}\chi_{j}\chi_{k}-(\chi_{i}\nabla_{j}\chi_{k}+\chi_{j}\nabla_{k}\chi_{i}+\chi_{k}\nabla_{i}\chi_{j})$
With the expression (57) the conformal Killing equation is satisfied for any choice of the constants.
Now we prove the opposite statement: in the metric (6), if $\mathsf{A,B,C}$ are those in thrm.1, then the tensor (13) is conformal Killing. This is expressed by condition (59), that becomes (61) after substitutions. The latter equation is identically satisfied. $\square$
Data availability statement
No data are available because of the nature of the research.